# Navigating the Metrics Maze: Reconciling Score Magnitudes and Accuracies

### v1

@@ 5.2 Recommendations for MT Evaluation
Use at least one additional metric of a different type; for example, BLEURT, which is reference-based and uses a different architecture from Comet.
@@ 6 Related Work
Automated evaluation metrics in NLP and MT have been under scrutiny for long time. Hovy and Ravichandran (2003) raised early doubts about BLEU. Callison-Burch et al. (2006) pointed to failure modes of BLEU and suggested it be used in more narrow situations. Post (2018) identified a problem with conflicting implementations of BLEU and offered a unified solution. The broader field of computer science has been concerned with what is a meaningful acceptance threshold of a metric (Mori et al., 2018). The acceptance thresholds are usually established to trade off risks in types of errors (Shatnawi et al., 2010). Kelley and Preacher (2012), in a study on effect sizes in psychology, summarize that effect sizes should be scaled appropriately. Similarly, Plonsky and Oswald (2014) directly ask what effect size is sufficient and note that it depends on the variance and that all acceptance thresholds are arbitrary.
@@ 7 Conclusion
In this work, we investigated the interpretation of deltas from automatic machine translation metrics. Although metrics have different ranges of scores, what ultimately matters to the practitioner is how score deltas are grounded in human ability to perceive those differences, which we judge by pairwise system-level accuracy on a large collection of human judgments. We empirically determined thresholds for popular metrics to align them on accuracy and provide a tool333kocmitom.github.io/MT-Thresholds that relates metrics to each other. Finally, we showed the importance of using metric-delta accuracy over -values: the former is stable across testset sizes.
@@ Appendix A Metric Implementation Details
There are many automatic metrics that has been developed. In our study, we focus only on the most used or the best performing ones. Here is the description, reason for their selection and details of implementations used. For metric quality, please, refer to Section 5.1 or Kocmi et al. (2021); Freitag et al. (2023).
Out of the lexical-based metrics, we select three of them, which are the most used. However, we emphasize that these metrics should no longer be used for MT evaluation (Freitag et al., 2022). We use SacreBLEU (Post, 2018) in version 2.3.1 with default setting:
@@
BLEURT (Pu et al., 2021): the best performing Bleurt model
@@ Appendix B ToShip23 Dataset Details
For this work, we introduce and utilize an extended version of a non-public ToShip23 dataset. The main changes of a dataset are almost twice as many system pairs as in ToShip21 (Kocmi et al., 2021); more than ten new domains and new parallel testsets; improved human evaluation protocol; and evaluating the latest state-of-the-art MT models.
The parallel testsets for evaluating MT models that we use in the extended part are mostly a collection of non-published human translated sentences. We focus on using testsets in authentic direction, from original source into human translated reference (avoiding reverse testsets whenever possible, Toral et al., 2018). In contrast to ToShip21, which uses mainly two domains (news and speech), we extended the list of focus domains by more than ten.
Regarding the number of languages, we reduced the total number of languages in the ToShip23 from 101 to 94. The removed languages are those which are not supported by either BERT (Devlin et al., 2019) or XLM-RoBERTa (Conneau et al., 2020) – language models used in the most popular metrics – therefore, we couldn’t include those languages in our analysis.
The MT systems being part of the dataset are coming from the same distribution as in ToShip21, but evaluating the most recent state-of-the-art models including a limited number of LLM based translations.
Lastly, we improved the human evaluation protocol, moved from source-based DA towards DA+SQM (Kocmi et al., 2022).
Appendix C Number of System Pairs in a Bin
In our work, we fixed the number of systems in a bin for given metric delta to 300 system pairs. We now show how this decision affected our evaluation. To this end, we show various bin sizes in Figure 9. The bin width works as a smoothing parameter. With bin size of 100 system pairs, the curve fluctuates significantly, especially as one system pair transfer into 1% change on the accuracy scale.
We decided to set the parameter to 300 system pairs as that is already a bit smoother curve, while it is not too wide that the epsilon around the investigated delta would not be too high. However, this parameter should be re-investigated in the future works.
Appendix D Extending Plots to All Metrics

### v2

@@ 5.2 Recommendations for MT Evaluation
Use at least one additional metric of a different type; e. g. BLEURT20, which is reference-based and uses a different architecture from Comet.
@@
Do not use BLEU, ChrF, or spBLEU to evaluate unrelated systems.
In addition, employ caution when using the same metric for evaluation that was used during training, as this practice may lead to artificially inflated results. For instance, it is advisable not to evaluate with the same metric used for Minimum Bayes Risk Decoding (Freitag et al., 2022a), QE metric used for corpus filtering (Peter et al., 2023), or avoid using metrics built on the same model as the translation system because LLMs tend to favor outputs generated by themselves (Liu et al., 2023).
@@ 6 Related Work
Automated metrics in NLP and MT have been under scrutiny for a long time. Hovy and Ravichandran (2003) raised early doubts about BLEU. Callison-Burch et al. (2006) pointed to failure modes of BLEU and suggested it be used in more narrow situations. Post (2018) identified a problem with conflicting implementations of BLEU and offered a unified solution. The broader field of computer science has been concerned with what is a meaningful acceptance threshold of a metric (Mori et al., 2018). The acceptance thresholds are usually established to trade off risks in types of errors (Shatnawi et al., 2010). Kelley and Preacher (2012), studying effect sizes in psychology, summarize that effect sizes should be scaled appropriately. Alike, Plonsky and Oswald (2014) ask what effect size suffices and note its dependence on the variance and that all acceptance thresholds are arbitrary.
@@ 7 Conclusion
In this work, we investigated the interpretation of deltas from automatic machine translation metrics. Although metrics have different ranges of scores, what ultimately matters to the practitioner is how score deltas are grounded in human ability to perceive those differences, which we judge by pairwise system-level accuracy on a large collection of human judgments. We empirically determined thresholds for popular metrics to align them on accuracy and provide a tool00footnotemark: 0 that relates metrics to each other. Finally, we showed the importance of using metric-delta accuracy over -values: the former is stable across testset sizes.
@@
Almost all MT systems used in this meta-evaluation are not based on LLMs. Therefore, we may observe different behaviour of automatic metrics when evaluating LLM-based models.
@@
Ethics Statement
The human annotators have been compensated considerably higher than the minimum wage standards in their respective countries. This commitment reflects our dedication to fair labor practices and the well-being of those contributing to our work.
@@ Appendix A Metric Implementation Details
There are many automatic metrics that has been developed. In our study, our selection of metrics has focus on either the most used metrics or the best performing ones. Here is the description, reason for their selection and details of implementations used. For metric quality, please, refer to Section 5.1 or Kocmi et al. (2021); Freitag et al. (2023). However, we plan to extend the list of automatic metric even after paper publishing. For other metrics and models, see https://github.com/kocmitom/MT-Thresholds/.
We are aware that the list is heavily affected by COMET variant models. However, when investigating best performing metrics from Freitag et al. (2023), we can see that most are either based on COMET framework, they are not publicly available, or build on propritetary LLMs.
Out of the lexical-based metrics, we select three of them, which are the most used. However, we emphasize that these metrics should no longer be used for MT evaluation (Freitag et al., 2022b). We use SacreBLEU (Post, 2018) in version 2.3.1 with default setting:
@@
BLEURT20 (Pu et al., 2021): the best performing Bleurt model
@@ Appendix B ToShip23 Dataset Details
For this work, we introduce and analyze an extended version of a non-public ToShip23 dataset. The main changes of a dataset are almost twice as many system pairs as in ToShip21 (Kocmi et al., 2021); more than ten new domains and new parallel testsets; improved human evaluation protocol; and evaluating the latest state-of-the-art MT models.
The parallel testsets for evaluating MT models that we use in the extended part are mostly a collection of non-published human translated sentences. We focus on using testsets in authentic direction, from original source into human translated reference (avoiding reverse testsets whenever possible, Toral et al., 2018). In contrast to ToShip21, which uses mainly two domains (news and speech), we extended the domains by more than ten.
We reduced the total number of languages in the ToShip23 from 101 to 94. The removed languages are those which are not supported by either BERT (Devlin et al., 2019) or XLM-RoBERTa (Conneau et al., 2020) – language models used in the most popular metrics – therefore, we could not include those languages in our analysis.
The MT systems being part of the dataset are coming from the same distribution as in ToShip21, but evaluating the most recent state-of-the-art models including a limited number of LLM based translations. Lastly, we improved the human evaluation protocol, moved from source-based DA towards DA+SQM (Kocmi et al., 2022).
Appendix C Metrics Disagreement on Ranking
Automatic metrics often disagree on a ranking which system is better even for large enough deltas. We illustrate this phenomena in this section.
We use the mostly unwritten (and long-debunked (Mathur et al., 2020a)) operating assumption that +1–2 BLEU points denotes a significant finding as an anchor point to illustrate the range of metric deltas on a subset of systems in Figure 8. This figure reports metric deltas for six randomly-selected system pairs from WMT23 data, whose delta was roughly 1 BLEU.
As we can see in Figure 8, while for first two system pairs, all metrics and humans agree on the system ranking, it is not the case for later four system pairs. For example, even Comet score of 3.4 (fifth system pair) may result in disagreement with humans.
Appendix D Number of System Pairs in a Bin
In our work, we fixed the number of systems in a bin for given metric delta to 300 system pairs. We now show how this decision affected our evaluation. To this end, we show various bin sizes in Figure 9. The bin width works as a smoothing parameter. With bin size of 100 system pairs, the curve fluctuates, especially as one system pair transfer into 1% change on the accuracy scale.
We set the parameter to 300 system pairs because that is already a smoother curve, while not too wide so that the epsilon around the investigated delta is also not too high. However, this parameter should be re-investigated in the future works.